# Why $w\neq-1$? Anthropic Selection in a $\Lambda$ + Axion Dark Energy Model

### v2

@@ Abstract
We study a dark energy model composed of a bare negative cosmological constant and a single ultra-light axion, motivated by the string axiverse. Assuming that intelligent observers arise and observe, as in our universe, the onset of dark-energy-driven acceleration following matter domination, and that this acceleration persists to the present, we derive nontrivial constraints on both the axion mass and the bare cosmological constant. The axion mass is bounded from above to avoid fine-tuning of the initial misalignment angle near the hilltop, and from below because too light axions cannot achieve accelerated expansion due to their limited energy budget. As a result, the anthropically allowed axion mass range typically lies around $m=\mathcal{O}(10)\,H_{0}$ for a decay constant close to the Planck scale, where $H_{0}$ is the observed value of the Hubble constant. In this framework, the dark energy equation-of-state parameter $w_{0}$ generically deviates from −$-1$ by $\mathcal{O}(0.1)$, providing a natural explanation for why $w\neq-1$ may be expected. We also find that, for a decay constant slightly smaller than the Planck scale, the peak value of dark energy density is significantly smaller than the anthropic bound on the cosmological constant and can be close to the observed value. These outcomes are intriguingly consistent with recent DESI hints of time-varying dark energy, and offer a compelling anthropic explanation within the $\Lambda$ + axion framework.
††preprint: TU-1263
@@ I Introduction
The accelerated expansion of the universe is now well established through various cosmological observations, including Type Ia supernovae Perlmutter et al. (1999); Riess et al. (1998), the cosmic microwave background Aghanim et al. (2020), and large-scale structure Alam et al. (2017). This phenomenon is commonly attributed to dark energy, which constitutes roughly 70% of the current energy density of the universe. Despite its dominant role, the physical origin of dark energy remains one of the most profound mysteries in modern cosmology. A wide variety of theoretical models have been proposed to explain it, ranging from modifications of gravity Carroll et al. (2004) to dynamical scalar fields Ratra and Peebles (1988); Caldwell et al. (1998). Nevertheless, the simplest and most conservative candidate is a cosmological constant Weinberg (1989).
@@
Among the proposed solutions to the cosmological constant problem, the anthropic principle Weinberg (1987, 1989) has gained significant attention, especially in the context of the string landscape Susskind (2003); Polchinski (2006); Ellis and Smolin (2009). If the landscape contains a vast number of vacua with different vacuum energies, it is natural to expect that observers can only exist in regions of the multiverse where the cosmological constant falls within a narrow anthropic window: too large a value would suppress galaxy formation, while a negative value would lead to rapid recollapse. This line of reasoning offers a statistical explanation for the observed smallness of the cosmological constant.111 In fact, the observed amount of dark energy is smaller than the anthropic upper bound by about three orders of magnitude, which may call for an explanation Sorini et al. (2024). See Appendix A for a possible resolution within our framework.
In this work, we go beyond the single-component picture in the anthropic argument and consider a scenario in which dark energy arises from the sum of a bare negative cosmological constant and the energy of an ultra-light axion. [^2] A related anthropic approach has been taken in the context of the strong CP problem, where the sum of a cosmological constant and an axion potential was considered as the effective cosmological constant Takahashi (2009); Kaloper (2019). Such ultra-light axions are generically predicted in string theory, where a large number of axions with masses spanning many orders of magnitude are expected—a scenario often referred to as the axiverse Arvanitaki et al. (2010). If the axion is sufficiently light, it can remain displaced from its potential minimum over cosmological timescales and contribute to the present-day dark energy density Svrcek (2006). The explanation of dark energy with a negative cosmological constant and an ultra-light axion has been explored in Refs. Ruchika et al. (2023); Luu et al. (2025). See also Refs. Cardenas et al. (2003); Dutta et al. (2020); Visinelli et al. (2019); Calderón et al. (2021); Sen et al. (2022); Adil et al. (2023); Menci et al. (2024) for related works.
@@
Remarkably, within this constrained parameter space, the dark energy equation-of-state parameter $w$ is generically predicted to deviate from −$-1$ by $\mathcal{O}(0.1)$. This deviation arises naturally from the anthropic selection and is consistent with recent DESI results hinting at time-varying dark energy Adame et al. (2025); Abdul Karim et al. (2025a, b), providing a compelling explanation within the $\Lambda$ + axion framework. See Refs. Tada and Terada (2024); Yin (2024); Cortês and Liddle (2024); Bhattacharya et al. (2024a); Mukherjee and Sen (2024); Notari et al. (2024); Jia et al. (2025); Hernández-Almada et al. (2024); Bhattacharya et al. (2024b); Berbig (2025); Wolf et al. (2025); Chakraborty et al. (2025); Borghetto et al. (2025); Nakagawa et al. (2025); Lee et al. (2025); Colgáin et al. (2025); Santos et al. (2025) for various explanations of the DESI results.
@@ II $\Lambda$ + Axion dark energy model
As a concrete realization of the anthropic scenario outlined in the introduction, we consider a model in which dark energy is composed of a negative cosmological constant, $\rho_{\Lambda}<0$, and a single axion field $\phi$. While the string axiverse generically contains multiple axions, we focus on a single axion field for simplicity. This approximation is justified if the axion mass spectrum is hierarchical: heavier axions would have already settled into their minima, while much lighter ones remain effectively frozen, and their energy densities can be effectively absorbed into $\rho_{\Lambda}$.
@@
where $m$ is the axion mass, $f$ is its decay constant, and we choose the origin, $\phi=0$, to be the maximum of the potential. We set $f=M_{\mathrm{Pl}}$ throughout the main text, and we will return to the case with smaller decay constants in a later discussion. We also assume that the axion field is spatially homogeneous.
@@
where $\rho_{\mathrm{DE,\mathrm{i}}}$ is the initial energy density of dark energy. We refer to this as a conditional anthropic bound. The lower bound reflects the need for the axion to uplift the negative vacuum energy and initially behave like a cosmological constant in order to drive accelerated expansion. The upper bound ensures that galaxy formation occurs before the axion starts oscillating. Note that the upper bound, $\rho_{\Lambda}^{\mathrm{max}}$, is derived under the assumption that the dark energy component solely consists of a cosmological constant. In our scenario, however, the axion should remain around the hilltop for accelerated expansion to persist until now, and so the dark energy density is expected to evolve only slowly. Therefore, we adopt $\rho_{\Lambda}^{\mathrm{max}}$ as the effective upper bound on the initial value of the total dark energy density.
@@
$\phi(t)$ $=\phi_{\mathrm{i}}\frac{\sinh(\sqrt{1+A}\tilde{t})}{\sqrt{1+A}\sinh(\tilde{t})}\ ,$ (13)
@@
×$\times\left[\frac{\cosh(\sqrt{1+A}\tilde{t}_{0})}{\sinh(\tilde{t}_{0})}-\frac{\cosh(\tilde{t}_{0})\cosh(\sqrt{1+A}\tilde{t}_{0})}{\sqrt{1+A}\sinh^{2}(\tilde{t}_{0})}\right]\ ,$ (16)
@@
We evaluate the evolution of the axion and $w_{0}$ both numerically and analytically. In Fig. 1, we show the parameter region allowed by the conditional anthropic bound for $m=100H_{0}$, $10H_{0}$, and $H_{0}$. We fix the range of $\rho_{\Lambda}$ following the condition (7). Note that the scales of the horizontal and vertical axes are different in the three panels. In the gray-shaded regions, our requirements are not satisfied: the universe never experiences accelerated expansion, or the accelerated expansion does not last until the current time. We also show $w_{0}$ by the color shading for the allowed regions. Smaller values of $\theta_{\mathrm{i}}$ lead to smaller $w_{0}$, as the axion remains closer to the hilltop. The required tuning of $\theta_{\mathrm{i}}$ becomes more severe for small $\rho_{\Lambda}$ (i.e., large $|\rho_{\Lambda}|$) because the Hubble parameter is small, which allows the axion to roll down the potential more easily. For $m=H_{0}$, $w_{0}$ cannot deviate significantly from −$-1$, namely, the axion should remain around the hilltop at the current time, since otherwise the accelerated universe cannot be achieved due to the small energy budget of the axion.
@@
With the flat priors for $\theta_{\mathrm{i}}$ and $\rho_{\Lambda}$, we can evaluate the likelihood of $\ln m$ by integrating the area of the colored regions in Fig. 1 over a denser sampling of the mass. We show the likelihood of $\ln m$ in Fig. 2. The vertical axis is normalized so that the maximum value equals unity. We find a peak at $m=m_{\mathrm{peak}}\simeq 49.2H_{0}$, which is determined by $2m^{2}f^{2}=\rho_{\Lambda}^{\mathrm{max}}$. This behavior can be understood as follows. Assuming that the axion stays near the hilltop, the allowed range of $\rho_{\Lambda}$ is limited as −$-2m^{2}f^{2}+\rho_{m,0}/2\lesssim\rho_{\Lambda}\lesssim{\rm Min}[0,\rho_{\Lambda}^{\rm max}-2m^{2}f^{2}]$. For small $m$ satisfying $\rho_{m,0}/2\ll 2m^{2}f^{2}<\rho_{\Lambda}^{\rm max}$, the interval of $\rho_{\Lambda}$ is approximately proportional to $m^{2}$. So, the likelihood increases roughly ∝$\propto m^{2}$. For large $m$ satisfying $2m^{2}f^{2}>\rho_{\Lambda}^{\rm max}$, while the interval of the allowed $\rho_{\Lambda}$ becomes approximately equal to $\rho_{\mathrm{\Lambda}}^{\mathrm{max}}$, independent of $m$, the required tuning of $\theta_{\mathrm{i}}$ near the hilltop becomes exponentially severe, suppressing the likelihood. Note that the peak position of the axion mass distribution depends on $\rho_{\Lambda}^{\mathrm{max}}$ as $m_{\mathrm{peak}}^{2}=\rho_{\Lambda}^{\mathrm{max}}/(2f^{2})$. In particular, if there exists some unknown anthropic condition that imposes a much tighter upper bound on $\rho_{\Lambda}$, which is close to the observed dark energy density, the peak of likelihood comes to $m_{\mathrm{peak}}=\mathcal{O}(H_{0})$. Moreover, this estimate for $m_{\mathrm{peak}}$ is further corrected for smaller $f$. See Appendix A for the different choices of $\rho_{\Lambda}^{\mathrm{max}}$ and $f$.
@@
In Fig. 3, we show the probability distribution of $w_{0}$ for fixed $m$. Here, we sample $w_{0}$ from −$-1$ to −$-0.3$ in increments of $0.05$. Note that the accelerated expansion at the current time implies $w_{0}<-1/3$. We evaluate the probability by integrating the area with corresponding values of $w_{0}$ in Fig. 1 and normalize it such that the total equals unity. We find that while the most likely value of $w_{0}$ lies close to −$-1$, the deviation by ${\cal O}(0.1)$ is statistically significant especially for $m=100H_{0}$ and $10H_{0}$. Given that the mass likelihood of $\ln m$ has a peak at $10H_{0}<m<100H_{0}$, our results suggest that, in a universe where intelligent observers can observe the accelerated expansion, the dark energy equation-of-state parameter is likely to deviate from −$-1$ at the level of ${\cal O}(0.1)$.
@@
Finally, we show the probability distribution of $\rho_{\mathrm{DE,i}}$ with the flat priors for $\theta_{\mathrm{i}}$, $\rho_{\Lambda}$, and $\ln m$ in Fig. 4. In a similar way to Fig. 3, we integrate the area of the colored regions with $\rho_{\mathrm{DE,i}}$ in each bin for different values of $\ln m$ and sum the area over $\ln m$ sampled with a uniform interval. Here, we take the bin for $24000<\rho_{\mathrm{DE,i}}/\rho_{\mathrm{DE,obs}}<$ 2400 in increment of $100$. Note that the rightmost bin corresponds to $23062300<\rho_{\mathrm{DE,i}}/\rho_{\mathrm{DE,obs}}<$ 2306 due to the requirement (7) and thus has a lower probability. The probability distribution for $\rho_{\mathrm{DE,i}}$ spreads out over the region allowed by Eq. (7), and thus we cannot account for the hierarchy between $\rho_{\mathrm{DE,obs}}$ and $\rho_{\Lambda}^{\mathrm{max}}$ in this case. On the other hand, as we discuss in Appendix A, $\rho_{\mathrm{DE,i}}$ close to the observed value is favored when we choose a smaller $f$ such as $f=0.1M_{\mathrm{Pl}}$. Thus, for $f={\cal O}(0.1)M_{\mathrm{Pl}}$, $w_{0}$ naturally deviates from −$-1$ by ${\cal O}(0.1)$, and the dark energy density approaches the observed value, resolving the hierarchy between $\rho_{\rm DE,0}$ and $\rho_{\Lambda}^{\rm max}$.

### v3

@@ Abstract
We study a dark energy model composed of a bare negative cosmological constant and a single ultra-light axion, motivated by the string axiverse. Assuming that intelligent observers arise and observe, as in our universe, the onset of dark-energy-driven acceleration following matter domination, and that this acceleration persists to the present, we derive nontrivial constraints on both the axion mass and the bare cosmological constant. The axion mass is bounded from above to avoid fine-tuning of the initial misalignment angle near the hilltop, and from below because too light axions cannot achieve accelerated expansion due to their limited energy budget. As a result, the anthropically allowed axion mass range typically lies around $m=\mathcal{O}(10)\,H_{0}$ for a decay constant close to the Planck scale, where $H_{0}$ is the observed value of the Hubble constant. In this framework, the dark energy equation-of-state parameter $w_{0}$ generically deviates from 1-1 by $\mathcal{O}(0.1)$, providing a natural explanation for why $w\neq-1$ may be expected. We also find that, for a decay constant slightly smaller than the Planck scale, the peak value of dark energy density is significantly smaller than the anthropic bound on the cosmological constant and can be close to the observed value. These outcomes are intriguingly consistent with recent DESI hints of time-varying dark energy, and offer a compelling anthropic explanation within the $\Lambda$ + axion framework.
preprint: TU-1263
@@ I Introduction
The accelerated expansion of the universe is now well established through various cosmological observations, including Type Ia supernovae Perlmutter et al. (1999); Riess et al. (1998), the cosmic microwave background (CMB) Aghanim et al. (2020), and large-scale structure Alam et al. (2017). This phenomenon is commonly attributed to dark energy, which constitutes roughly 70% of the current energy density of the universe. Despite its dominant role, the physical origin of dark energy remains one of the most profound mysteries in modern cosmology. A wide variety of theoretical models have been proposed to explain it, ranging from modifications of gravity Carroll et al. (2004) to dynamical scalar fields Ratra and Peebles (1988); Caldwell et al. (1998). Nevertheless, the simplest and most conservative candidate is a cosmological constant Weinberg (1989).
@@
Among the proposed solutions to the cosmological constant problem, the anthropic principle Weinberg (1987, 1989) has gained significant attention, especially in the context of the string landscape Susskind (2003); Polchinski (2006); Ellis and Smolin (2009). If the landscape contains a vast number of vacua with different vacuum energies, it is natural to expect that observers can only exist in regions of the multiverse where the cosmological constant falls within a narrow anthropic window: too large a value would suppress galaxy formation, while a negative value would lead to rapid recollapse. This line of reasoning offers a statistical explanation for the observed smallness of the cosmological constant. [^1] In fact, the observed amount of dark energy is smaller than the anthropic upper bound by about three orders of magnitude, which may call for an explanation Sorini et al. (2024). See Appendix A for a possible resolution within our framework.
In this work, we go beyond the single-component picture in the anthropic argument and consider a scenario in which dark energy arises from the sum of a bare negative cosmological constant and the energy of an ultra-light axion. [^2] A related anthropic approach has been taken in the context of the strong CP problem, where the sum of a cosmological constant and an axion potential was considered as the effective cosmological constant Takahashi (2009); Kaloper (2019). Such ultra-light axions are generically predicted in string theory, where a large number of axions with masses spanning many orders of magnitude are expected—a scenario often referred to as the axiverse Arvanitaki et al. (2010). If the axion is sufficiently light, it can remain displaced from its potential minimum over cosmological timescales and contribute to the present-day dark energy density Svrcek (2006). The explanation of dark energy with a negative cosmological constant and an ultra-light axion has been explored in Refs. Ruchika et al. (2023); Luu et al. (2025a). See also Refs. Cardenas et al. (2003); Dutta et al. (2020); Visinelli et al. (2019); Calderón et al. (2021); Sen et al. (2022); Adil et al. (2023); Menci et al. (2024) for related works.
@@
Remarkably, within this constrained parameter space, the dark energy equation-of-state parameter $w$ is generically predicted to deviate from 1-1 by $\mathcal{O}(0.1)$. This deviation arises naturally from the anthropic selection and is consistent with recent DESI results hinting at time-varying dark energy Adame et al. (2025); Abdul Karim et al. (2025a, b), providing a compelling explanation within the $\Lambda$ + axion framework. See Refs. Tada and Terada (2024); Yin (2024); Cortês and Liddle (2024); Bhattacharya et al. (2024a); Mukherjee and Sen (2024); Notari et al. (2024); Jia et al. (2025); Hernández-Almada et al. (2024); Bhattacharya et al. (2024b); Berbig (2025); Wolf et al. (2025); Chakraborty et al. (2025); Borghetto et al. (2025); Nakagawa et al. (2025); Lee et al. (2025); Colgáin et al. (2025); Santos et al. (2025) for various explanations of the DESI results.
@@ II $\Lambda$ + Axion dark energy model
As a concrete realization of the anthropic scenario outlined in the introduction, we consider a model in which dark energy is composed of a negative cosmological constant, $\rho_{\Lambda}<0$, and a single axion field $\phi$. Here, the presence of a negative cosmological constant is an assumption motivated by string theory Kachru et al. (2003); Bousso and Polchinski (2000); Obied et al. (2018); Danielsson and Van Riet (2018); Demirtas et al. (2022), while $\rho_{\Lambda}<0$ is preferred from the observational data in the framework of $\Lambda$ + axion dark energy Luu et al. (2025a). While the string axiverse generically contains multiple axions, we focus on a single axion field for simplicity. This approximation is justified if the axion mass spectrum is hierarchical: heavier axions would have already settled into their minima, while much lighter ones remain effectively frozen, and their energy densities can be effectively absorbed into $\rho_{\Lambda}$.
@@
where $m$ is the axion mass, $f$ is its decay constant, and we choose the origin, $\phi=0$, to be the maximum of the potential. We set $f=M_{\mathrm{Pl}}$ throughout the main text, and we will consider the case with smaller decay constants in a later discussion. We also assume that the axion field is spatially homogeneous.
@@
where $\rho_{\mathrm{DE,\mathrm{i}}}$ is the initial energy density of dark energy. We refer to this as a conditional anthropic bound. The lower bound reflects the need for the axion to uplift the negative vacuum energy and initially behave like a cosmological constant in order to drive accelerated expansion. The upper bound ensures that galaxy formation occurs before the axion starts oscillating. Note that the upper bound, $\rho_{\Lambda}^{\mathrm{max}}$, is derived under the assumption that the dark energy component solely consists of a cosmological constant. In our scenario, however, the axion should remain around the hilltop for accelerated expansion to persist until now, and so the dark energy density is expected to evolve only slowly. Therefore, we adopt $\rho_{\Lambda}^{\mathrm{max}}$ as the effective upper bound on the initial value of the total dark energy density. Since the axion remains frozen in the early universe, we can set the initial time arbitrarily as long as it is well before the axion starts to evolve. Due to the upper bound on $\rho_{\mathrm{DE,\mathrm{i}}}$, the dark energy is negligible until well after the matter-radiation equality, and it does not affect big bang nucleosynthesis and recombination.
@@
$\phi(t)$ $=\phi_{\mathrm{i}}\frac{\sinh(\sqrt{1+A}\tilde{t})}{\sqrt{1+A}\sinh(\tilde{t})}\$, (13)
@@
$\times\left[\frac{\cosh(\sqrt{1+A}\tilde{t}_{0})}{\sinh(\tilde{t}_{0})}-\frac{\cosh(\tilde{t}_{0})\cosh(\sqrt{1+A}\tilde{t}_{0})}{\sqrt{1+A}\sinh^{2}(\tilde{t}_{0})}\right]\ ,$ (16)
@@
We evaluate the evolution of the axion and $w_{0}$ both numerically and analytically. In Fig. 1, we show the parameter region allowed by the conditional anthropic bound for $m=100H_{0}$, $10H_{0}$, and $H_{0}$. We fix the range of $\rho_{\Lambda}$ following the condition (7). Note that the scales of the horizontal and vertical axes are different in the three panels. In the gray-shaded regions, our requirements are not satisfied: the universe never experiences accelerated expansion, or the accelerated expansion does not last until the current time. We also show $w_{0}$ by the color shading for the allowed regions. Smaller values of $\theta_{\mathrm{i}}$ lead to smaller $w_{0}$, as the axion remains closer to the hilltop. The required tuning of $\theta_{\mathrm{i}}$ becomes more severe for small $\rho_{\Lambda}$ (i.e., large $|\rho_{\Lambda}|$) because the Hubble parameter is small, which allows the axion to roll down the potential more easily. For $m=H_{0}$, $w_{0}$ cannot deviate significantly from 1-1, namely, the axion should remain around the hilltop at the current time, since otherwise the accelerated universe cannot be achieved due to the small energy budget of the axion.
@@
With the flat priors for $\theta_{\mathrm{i}}$ and $\rho_{\Lambda}$, we can evaluate the likelihood of $\ln m$ by integrating the area of the colored regions in Fig. 1 over a denser sampling of the mass. We show the likelihood of $\ln m$ in Fig. 2. The vertical axis is normalized so that the maximum value equals unity. We find a peak at $m=m_{\mathrm{peak}}\simeq 49.2H_{0}$, which is determined by $2m^{2}f^{2}=\rho_{\Lambda}^{\mathrm{max}}$. This behavior can be understood as follows. Assuming that the axion stays near the hilltop, the allowed range of $\rho_{\Lambda}$ is limited as $2m^{2}f^{2}+\rho_{m,0}/2\lesssim\rho_{\Lambda}\lesssim{\rm Min}[0,\rho_{\Lambda}^{\rm max}-2m^{2}f^{2}]$. For small $m$ satisfying $\rho_{m,0}/2\ll 2m^{2}f^{2}<\rho_{\Lambda}^{\rm max}$, the interval of $\rho_{\Lambda}$ is approximately proportional to $m^{2}$. So, the likelihood increases roughly $\propto m^{2}$. For large $m$ satisfying $2m^{2}f^{2}>\rho_{\Lambda}^{\rm max}$, while the interval of the allowed $\rho_{\Lambda}$ becomes approximately equal to $\rho_{\mathrm{\Lambda}}^{\mathrm{max}}$, independent of $m$, the required tuning of $\theta_{\mathrm{i}}$ near the hilltop becomes exponentially severe, suppressing the likelihood. Note that the peak position of the axion mass distribution depends on $\rho_{\Lambda}^{\mathrm{max}}$ as $m_{\mathrm{peak}}^{2}=\rho_{\Lambda}^{\mathrm{max}}/(2f^{2})$. In particular, if there exists some unknown anthropic condition that imposes a much tighter upper bound on $\rho_{\Lambda}$, which is close to the observed dark energy density, the peak of likelihood comes to $m_{\mathrm{peak}}=\mathcal{O}(H_{0})$. Moreover, this estimate for $m_{\mathrm{peak}}$ is further corrected for smaller $f$. See Appendix A for the different choices of $\rho_{\Lambda}^{\mathrm{max}}$ and $f$.
@@
In Fig. 3, we show the probability distribution of $w_{0}$ for fixed $m$. Here, we sample $w_{0}$ from 1-1 to 0.3-0.3 in increments of $0.05$. Note that the accelerated expansion at the current time implies $w_{0}<-1/3$. We evaluate the probability by integrating the area with corresponding values of $w_{0}$ in Fig. 1 and normalize it such that the total equals unity. We find that while the most likely value of $w_{0}$ lies close to 1-1, the deviation by ${\cal O}(0.1)$ is statistically significant especially for $m=100H_{0}$ and $10H_{0}$. Given that the mass likelihood of $\ln m$ has a peak at $10H_{0}<m<100H_{0}$, our results suggest that, in a universe where intelligent observers can observe the accelerated expansion, the dark energy equation-of-state parameter is likely to deviate from 1-1 at the level of ${\cal O}(0.1)$.
@@
Finally, we show the probability distribution of $\rho_{\mathrm{DE,i}}$ with the flat priors for $\theta_{\mathrm{i}}$, $\rho_{\Lambda}$, and $\ln m$ in Fig. 4. In a similar way to Fig. 3, we integrate the area of the colored regions with $\rho_{\mathrm{DE,i}}$ in each bin for different values of $\ln m$ and sum the area over $\ln m$ sampled with a uniform interval. Here, we take the bin for $0<\rho_{\mathrm{DE,i}}/\rho_{\mathrm{DE,obs}}<2400$ in increment of $100$. Note that the rightmost bin corresponds to $2300<\rho_{\mathrm{DE,i}}/\rho_{\mathrm{DE,obs}}<2306$ due to the requirement (7) and thus has a lower probability. The probability distribution for $\rho_{\mathrm{DE,i}}$ spreads out over the region allowed by Eq. (7), and thus we cannot account for the hierarchy between $\rho_{\mathrm{DE,obs}}$ and $\rho_{\Lambda}^{\mathrm{max}}$ in this case. On the other hand, as we discuss in Appendix A, $\rho_{\mathrm{DE,i}}$ close to the observed value is favored when we choose a smaller $f$ such as $f=0.1M_{\mathrm{Pl}}$. Thus, for $f={\cal O}(0.1)M_{\mathrm{Pl}}$, $w_{0}$ naturally deviates from 1-1 by ${\cal O}(0.1)$, and the dark energy density approaches the observed value, resolving the hierarchy between $\rho_{\rm DE,0}$ and $\rho_{\Lambda}^{\rm max}$.